# Assessing (im)balance in signed brain networks

### v1

@@ IX Discussion
The study of the role played by negative connections within and between brain areas is a topic that has recently gained attention within the neuroscientific literature. The method that we have proposed here produces connections connotated by a negative sign as the result of a validation procedure that counts the number of times the corresponding series are found to be discordant [^22] Aggregating positive and negative contributions can, in fact, obscure the presence of anti-correlations. For instance, while the DMN and the DA network are connected by, approximately, $3\cdot 10^{5}$ positive links, they are also connected by, roughly, $9\cdot 10^{4}$ negative links: summing these numbers together would yield a net positive count, misleadingly hiding the presence of anti-correlated activity. and compare it with the prediction of a properly defined benchmark. Such a validation scheme allows us to answer a number of questions of neuroscientific interest.
@@
The first one concerns the frustration of the brain, at both the microscopic and the mesoscopic level. For what concerns the first one, the evidence that the median IBI value is positive - irrespectively from the benchmark employed to project our set of multivariate time series - lets us conclude that brain networks are indeed populated by a vast majority of (++)(+-+)( + - + ) triples - the ones reading ()(---)( - - - ) being almost absent. Interestingly, although the negative subgraph is more prone to inter-subject variability and less homogeneous than the positive one, the negative connections constituting the frustrated triples are mostly found between the groups of nodes constituting the DMN and the groups of nodes constituting task-positive regions such as the DA, the SM, the SVM and the VIS. As already mentioned, these results match the ones in Fox et al. (2005) as well as those discussed in Demertzi et al. (2022); Greicius et al. (2003); Hampson et al. (2010).
@@
Some contributions observe that the empirical level of (im)balance is smaller than the one characterising randomised configurations [^23] It should be, however, noticed that the results obtained in Saberi et al. (2021) depend on the chosen threshold, as also explicitly recognised by the authors. and deduce that the brain tends to stabilise itself Saberi et al. (2021). Our results, however, lead to a different conclusion: although it is true that the bSCM returns (slightly) less balanced configurations than the naïve ones, brain networks do not seem to minimise frustration: as confirmed by the distribution of $F$, the values assumed by it across the partitions characterising our and Yeo’s projections are quite larger than zero.
@@
It is, thus, of interest asking which principle is obeyed by the brain in order to self-organise its functioning. An answer seems to be provided by fig. 14: task-based partitions align with those recovered by minimising BIC; in a sense, thus, some sort of optimisation principle seems to be at work at the mesoscopic scale. Still, individuating the partition ‘best representing’ a signed brain network remains a challenging task, as different metrics provide different answers; nevertheless, a message whose validity appear as quite general seems to emerge: brain networks align with the (statistical variant) of the Relaxed Balanced Theory, as a vast majority of positive signs populates off-diagonal blocks as well.
@@ X Conclusions
A critical, yet often overlooked, aspect of functional brain data is the role played by negative links: frequently ignored due to the interpretative challenges brought along, emerging evidence indicates that negative links play a fundamental role in shaping the (structural and dynamical) properties of resting-state networks, especially those related to self-stabilisation. As excluding negative links, thus, means underestimating the brain ability to function properly, we have proposed a validation technique allowing statistically significant signed links (be they positive or negative) to emerge out of the noise accompanying raw data. To this aim, two, different benchmarks have been proposed, filtering multivariate time series to a different extent and letting diverse patterns emerge out of the wealth of empirical observations: in particular, while the bSRGM emphasizes structural balance by levelling out the differences between time series, the bSCM preserves their heterogeneity, thus inducing sparser projections that retain more information about the original, bipartite structure. In particular, our analysis points out that brain networks are, indeed, frustrated, the patterns induced by signed connections overlapping to a quite large extent with those individuated by Yeo’s task-oriented partitions and returning the picture of a system aligning (better) with the statistical variant of the Relaxed Balance Theory.
A future research direction is that of employing our algorithm to carry out both a ‘weighted-to-binary’ and a ‘weighted-to-weighted’ analysis, i.e. obtain statistically validated binary and weighted projections of multivariate weighted time series, defined by the ‘rule’ that any two nodes sharing a number of (either concordant or discordant) relationships characterised by a significantly large weight are connected by an edge ‘dressed’ with the validated weight.
@@ Appendix A Probabilistic models for binary undirected bipartite signed networks
$C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining $M-1$ constraints represent proper, topological properties. Such an optimisation procedure defines the expression
@@ Appendix B Numerical optimisation of likelihood functions
In order to numerically solve the systems of equations defining the bSCM, we can follow the guidelines provided in Vallarano et al. (2021): more specifically, we will adapt the iterative recipe provided there to our setting. First, let us notice that, upon posing $x_{i}\equiv e^{-\alpha_{i}}$, $y_{i}\equiv e^{-\gamma_{i}}$, $\forall\>i$ and $z_{t}\equiv e^{-\delta_{t}}$, $w_{t}\equiv e^{-\eta_{t}}$, $\forall\>t$, the system of equations defining the bSCM can be re-written as
@@
in order for our iterative recipe to converge, an appropriate vector of initial conditions needs to be chosen; here, we have opted the following one: $x_{i}=k_{i}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>i$, $y_{i}=k_{i}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>i$, $z_{t}=\kappa_{t}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>t$ and $w_{t}=\kappa_{t}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>t$.
Moreover, we have adopted two, different stopping criteria: the first one is a condition on the Euclidean norm of the vector of differences between the values of the parameters at subsequent iterations, i.e. $||\Delta\vec{\theta}||_{2}=\sqrt{\sum_{i=1}^{N}(\Delta\theta_{i})^{2}}\leq 10^{-8}$; the second one is a condition on the maximum number of iterations of our iterative algorithm, set to $10^{3}$.
@@
Table 1, providing information on the basic statistics of 555, illustrative time series coming from the HCP-Young Adult dataset shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining the bSCM: our method is, overall, fast and accurate, as the absolute errors are around $10^{-3}$ and the relative errors are around $10^{-6}$, the time employed to achieve such an accuracy being always less than a minute.
Basic statistics of time series bSCM $N$ $T$ $B^{+}$ $B^{-}$ $c^{+}$ $c^{-}$ MAE MRE Time (s) HCP106016 116 2.400 142.407 135.993 0.512$\simeq$ 0.512≃ 0.512 0.488$\simeq$ 0.488≃ 0.488 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.94\cdot 10^{-7}$ 28.8$\simeq$ 28.8≃ 28.8 HCP857263 116 2.400 136.376 142.024 0.490$\simeq$ 0.490≃ 0.490 0.510$\simeq$ 0.510≃ 0.510 $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.90\cdot 10^{-7}$ 33.3$\simeq$ 33.3≃ 33.3 HCP125525 116 2.400 139.054 139.346 0.499$\simeq$ 0.499≃ 0.499 0.501$\simeq$ 0.501≃ 0.501 $\simeq 1.26\cdot 10^{-3}$ $\simeq 1.17\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP100408 116 2.400 137.804 140.596 0.495$\simeq$ 0.495≃ 0.495 0.505$\simeq$ 0.505≃ 0.505 $\simeq 1.30\cdot 10^{-3}$ $\simeq 1.15\cdot 10^{-6}$ 37.8$\simeq$ 37.8≃ 37.8 HCP130316 116 2.400 137.196 141.204 0.493$\simeq$ 0.493≃ 0.493 0.507$\simeq$ 0.507≃ 0.507 $\simeq 1.97\cdot 10^{-3}$ $\simeq 1.89\cdot 10^{-6}$ 37.7$\simeq$ 37.7≃ 37.7
Table 1: Performance of our fixed-point algorithm to solve the system of equations defining the bSCM on 555, illustrative time series coming from the HCP-Young Adult dataset (each pedex refers to a subject). $N$ is the number of brain regions, $T$ is the duration of the related time series, $B^{+}$ is the overall number of positive values, $B^{-}$ is the overall number of negative values, $c^{+}=B^{+}/(N\cdot T)$ is the density of positive values and $c^{-}=B^{-}/(N\cdot T)$ is the density of negative values. Our fixed-point algorithm leads to MAE values around $10^{-3}$ and MRE values of the order of $10^{-6}$ in less than forty seconds.
@@ Appendix C Degree distributions of cortical and subcortical areas
As anticipated in the main text, subcortical areas tend to have a larger negative degree than cortical areas. This is true in both a subject-wise (see fig. 16) and a pooled fashion (see fig. 17)
Figure 16: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark. Analyses have been carried out in a subject-wise fashion (i.e. for subjects 106016106016106016 and 857263857263857263).
Figure 17: Complementary cumulative distribution functions (CCDFs) of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark. Analyses have been carried out in a pooled fashion (i.e. across the 100100100 subjects considered for the present analysis).

### v3

@@ IX Discussion
Understanding the role played by negative connections within and between brain areas is a topic that has recently gained attention within the neuroscientific literature. The method proposed here produces connections characterized by a negative sign as the result of a validation procedure that counts the number of times the corresponding series are found to be discordant and compares it with the prediction of a properly defined benchmark.
@@
Such a validation scheme allows us to address several neuroscientific questions. The first concerns the frustration of the brain at both the microscopic and mesoscopic levels. For the first one, the evidence that the median IBI value is positive - regardless of the benchmark employed to project our set of multivariate time series - lets us conclude that brain networks are populated by a non-negligible number of (++)(+-+) and ()(---) triads. Interestingly, the negative connections constituting the frustrated triads are mostly found within the groups of subcortical and limbic areas and between the groups of nodes constituting the DMN and those constituting task-positive regions such as the DA, the SM, the SVM, and the VIS; as already mentioned, these results match the ones discussed in Greicius2003functional; Fox2005Human; Hampson2010Functional; Demertzi2022.
@@
Some contributions observe that the empirical level of (im)balance is smaller than the one characterising randomised configurations [^21] It should be, however, noticed that the results obtained in Saberi2021 depend on the chosen threshold, as explicitly recognised by the authors themselves. and deduce that the brain tends to stabilise itself Saberi2021. Our results, however, lead to a different conclusion, i.e. that brain networks do not seem to minimise frustration: on the contrary, the values assumed by $F$ across Yeo’s partitions and those characterising our projections are quite larger than zero.
It is, thus, of interest to wonder which principle is obeyed by the brain to self-organise its functioning. An answer seems to be provided by fig. 13: Yeo’s partitions align with those returned by BIC minimisation, an advantage of which is that of acting in a maximally ‘agnostic’ way when coming to gather signed connections - differently from modularity maximisation, or closely related approaches Zhan2017, that seek for the group-wise structure best aligning with the Traditional Balance Theory.
@@
In a sense, thus, some sort of optimisation principle seems to be at work at the mesoscopic scale. Still, individuating the partition ‘best representing’ a signed brain network remains a challenging task, as different metrics provide different answers. Nevertheless, a message whose validity appear as quite general emerges: brain networks align with the (statistical variant) of the Relaxed Balanced Theory, as negative (positive) signs can populate diagonal (off-diagonal) blocks.
For what concerns the physiological mechanisms behind the emergence of negative connections, instead, we redirect the interested reader to more technical contributions Harita2024.
@@ X Conclusions
A critical, yet often overlooked, aspect of functional brain data is the role played by negative links: frequently ignored due to the interpretative challenges brought along, emerging evidence indicates negative links to be fundamental in shaping resting-state networks - especially the properties affecting self-stabilisation. As excluding negative links, thus, means underestimating the brain ability to function properly, we have proposed a validation technique allowing statistically significant signed links to emerge out of the noise accompanying raw data; two, different benchmarks have been employed, filtering multivariate time series to a different extent and letting diverse patterns emerge out of the wealth of empirical observations: while the bSRGM levels out the differences between time series, the bSCM preserves their heterogeneity, thus inducing sparser projections that retain more information about the original structure. In particular, our analysis points out that brain networks are, indeed, frustrated, the patterns induced by signed connections overlapping (in some cases, to a quite large extent) with Yeo’s task-based partitions and returning the picture of a system aligning (better) with the statistical variant of the Relaxed Balance Theory.
A future research direction is that of employing our algorithm to obtain statistically validated binary and weighted projections of multivariate weighted time series, defined by the ‘rule’ that any two nodes involved in a number of relationships characterised by a significantly large weight are connected by an edge ‘dressed’ with such a validated weight.
@@ Appendix A Probabilistic models for binary undirected bipartite signed networks
$C_{0}\equiv\langle C_{0}\rangle\equiv 1$ sums up the normalisation condition and the remaining M1M-1 constraints represent proper, topological properties. Such an optimisation procedure defines the expression
@@ Appendix B Numerical optimisation of likelihood functions
In order to numerically solve the system of equations defining the bSCM, we can follow the guidelines provided in vallarano2021: more specifically, we will adapt the iterative recipe provided there to our setting. First, let us notice that, upon posing $x_{i}\equiv e^{-\alpha_{i}}$, $y_{i}\equiv e^{-\gamma_{i}}$, $\forall\>i$ and $z_{t}\equiv e^{-\delta_{t}}$, $w_{t}\equiv e^{-\eta_{t}}$, $\forall\>t$, the system of equations defining the bSCM can be re-written as
@@
in order for our iterative recipe to converge, an appropriate vector of initial conditions needs to be chosen; here, we have opted for the following one: $x_{i}=k_{i}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>i$, $y_{i}=k_{i}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>i$, $z_{t}=\kappa_{t}^{+}(\mathbf{B}^{*})/\sqrt{L^{+}(\mathbf{B}^{*})}$, $\forall\>t$ and $w_{t}=\kappa_{t}^{-}(\mathbf{B}^{*})/\sqrt{L^{-}(\mathbf{B}^{*})}$, $\forall\>t$.
Moreover, we have adopted two, different stopping criteria: the first one is a condition on the Euclidean norm of the vector of differences between the values of the parameters at subsequent iterations, i.e. $\Delta\vec{\theta}||_{2}=\sqrt{\sum_{i=1}^{N}(\Delta\theta_{i})^{2}}\leq 10^{-8}$; the second one is a condition on the maximum number of iterations of our iterative algorithm, set to $10^{3}$.
@@
Table 1, providing information on the basic statistics of $5$, illustrative time series coming from the HCP-Young Adult dataset shows the time employed by our algorithm to converge as well as its accuracy in reproducing the constraints defining the bSCM: our method is, overall, fast and accurate, as the absolute error is steadily around $10^{-3}$ and the relative error is steadily around $10^{-6}$, the time employed to achieve such an accuracy being always less than a minute.
Basic statistics of time series bSCM NN $T$ $^{+} BB^{-}$ $^{+} cc^{-}$ MAE MRE Time (s) HCP100408 116 2.400 137.804 140.596 $\simeq 0.495$ $\simeq 0.505$ $\simeq 1.30\cdot 10^{-3}$ $\simeq 1.15\cdot 10^{-6}$ $\simeq 37.8$ HCP106016 116 2.400 142.407 135.993 $\simeq 0.512$ $\simeq 0.488$ $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.94\cdot 10^{-7}$ $\simeq 28.8$ HCP125525 116 2.400 139.054 139.346 $\simeq 0.499$ $\simeq 0.501$ $\simeq 1.26\cdot 10^{-3}$ $\simeq 1.17\cdot 10^{-6}$ $\simeq 37.8$ HCP130316 116 2.400 137.196 141.204 $\simeq 0.493$ $\simeq 0.507$ $\simeq 1.97\cdot 10^{-3}$ $\simeq 1.89\cdot 10^{-6}$ $\simeq 37.7$ HCP857263 116 2.400 136.376 142.024 $\simeq 0.490$ $\simeq 0.510$ $\simeq 1.10\cdot 10^{-3}$ $\simeq 9.90\cdot 10^{-7}$ $\simeq 33.3$
Table 1: Performance of our fixed-point algorithm to solve the system of equations defining the bSCM on 55, illustrative time series coming from the HCP-Young Adult dataset (each pedex refers to a subject): NN is the number of brain regions, $T$ is the duration of the related time series, $B^{+}$ is the overall number of positive values, $B^{-}$ is the overall number of negative values, $c^{+}=B^{+}/(N\cdot T)$ is the density of positive values and $c^{-}=B^{-}/(N\cdot T)$ is the density of negative values. Our fixed-point algorithm leads to MAE values around $10^{-3}$ and MRE values of the order of $10^{-6}$ in less than forty seconds.
@@ Appendix C Degree distributions of cortical and subcortical areas
As anticipated in the main text, subcortical areas tend to have a larger negative degree than cortical areas.
Figure 16: Cumulative distributions of positive ($k^{+}$) and negative ($k^{-}$) degrees for cortical and subcortical areas, per benchmark. Analyses have been carried out in a pooled fashion (i.e. across the $100$ subjects considered for the present analysis).